# Information Accessibility Limits in Structured NP Search

### 05 May 2026 v2

@@ 3.3 Mechanism: Information Masking and Intrinsic Limitation
The rank-one construction further clarifies this point. Once a subset with minimal principal minor has been identified, a simple perturbation suffices to produce a violation. Thus, the difficulty does not lie in modifying the matrix, but in acquiring the information needed to locate such a subset.
The information-accessibility barrier is therefore intrinsic: even under idealized computation, the total information available is limited by what queries reveal, and the barrier persists independently of numerical precision.
In this sense, information is the fundamental resource: no algorithm can recover the solution without acquiring the necessary information, regardless of computational power or precision.
@@ 4 Statistical Indistinguishability and Information Accessibility
The information-accessibility barrier can be understood from the statistical behavior of the response transcript.
@@
Consequently, the response transcript is asymptotically independent of $W$. Since $\mathcal{F}_{q}$ is almost deterministic, it carries vanishing information about the identity of the violating subset. Formally,
@@
Therefore, polynomially many queries cannot identify the violating subset: specifying $W$ requires $\Theta(N)$ bits, while the response transcript reveals only a vanishing amount. This recovers the same conclusion as the Fano-based argument.
Interpretation. The argument above establishes that, under the oracle interaction model, transcript distributions corresponding to different instances are statistically indistinguishable except on exponentially rare events. In particular, the observable signal distinguishing different instances is extremely weak.
In the near-boundary regime, instances with a unique violating subset lie close to P-matrix instances, so perturbations affect observable quantities only slightly. This provides a structural explanation for the weak signal observed in the oracle model.
We emphasize that the indistinguishability result is derived under the oracle interaction model. However, the near-boundary construction suggests that the weakness of the observable signal may not be solely an artifact of this model, but may reflect an intrinsic property of the instances themselves: the information specifying the solution is globally encoded while only weakly expressed in observable quantities.
Conjecture. There exist distributions over instances of the form $A(u,v)$ under which finding some violating subset $\alpha$ with
@@
can be achieved with non-negligible probability, while recovering a specific violating subset $w^{\star}$ (for example, one achieving a prescribed value such as $\det(A(u,v)_{w^{\star}})\approx-\epsilon$) remains computationally hard on average.
This gap is not captured by the oracle lower bound established in this work, and reflects a stronger form of information inaccessibility: the solution is globally encoded while the resulting observations are statistically indistinguishable.
This conjecture suggests that the information-accessibility barrier may extend beyond oracle models, pointing toward a broader class of structured problems in which global encoding and limited observability jointly induce intrinsic computational difficulty.
@@ 5 Concluding Remarks
The near-boundary construction arises from intrinsic properties of P-matrices rather than the interaction model. The parametric representation drives the system toward the boundary of feasibility, where violations emerge while the underlying structure is preserved, providing a controlled setting in which information-accessibility barriers can be isolated and analyzed.
Compared to more complex problems such as $\mathrm{3\text{-}SAT}$, where multiple interacting factors contribute to computational difficulty, the P-matrix violation problem offers a more transparent setting for studying information-accessibility barriers. In this sense, the construction serves as a clean experimental model of search hardness.
Our analysis relies on fundamental limits on information acquisition, establishing impossibility results via mutual information and Fano’s inequality. Together with the near-boundary construction, this shows that the observed barrier is an inherent consequence within the considered interaction model.
A key limitation of the present analysis is the use of a uniform prior and a simplified access model. In realistic settings, such as 3-SAT, clause overlap and adaptive inference introduce complex dependencies that may increase the effective rate of information acquisition. This perspective thus offers a complementary lens to traditional complexity analysis, emphasizing the accessibility of information rather than computational resources alone.
Algorithms such as PPSZ [2] demonstrate that certain structural features can be exploited to improve search performance; however, their behavior also suggests that information extraction remains limited and must be accumulated over many steps. Understanding how structural properties enable higher rates of information extraction, and whether fundamental limits exist within specific interaction models, is an important direction for future research. More broadly, a central open problem is to characterize the regimes in which information-accessibility barriers arise and to quantify the limits on information flow imposed by problem structure.
A further direction is to study the information-acquisition dynamics of structure-exploiting algorithms such as PPSZ. One may conjecture that such algorithms exhibit distinct phases: an initial phase in which structural features (e.g., critical clauses) enable relatively effective information extraction, followed by a regime in which the remaining uncertainty becomes less accessible, leading to a reduced effective information-acquisition rate. Notably, this transition need not reflect a loss of structural constraints - such as the existence of critical clauses - but rather a reduction in their accessibility under the algorithmic interaction. It is natural to ask whether this behavior reflects an intrinsic property of structured search under the given interaction model. Characterizing this transition - empirically and theoretically - and understanding its dependence on instance structure and variable ordering are promising directions for future research.

### 06 May 2026 v3

@@ 3.3 Mechanism: Information Masking and Intrinsic Limitation
The rank-one construction clarifies that the difficulty lies not in creating a violation - a small perturbation suffices once an appropriate subset is known - but in acquiring the information needed to locate such a subset.
The information-accessibility barrier is therefore intrinsic: even under idealized computation, recovery is limited by the information revealed through queries rather than by computational power or numerical precision.
@@ 4 Statistical Indistinguishability and Information Accessibility
The information-accessibility barrier can be understood through the statistical behavior of the response transcript.
@@
Therefore, the response transcript is asymptotically independent of $W$. Since $\mathcal{F}_{q}$ is almost deterministic, it carries only vanishing information about the identity of the violating subset. Formally,
@@
Consequently, polynomially many queries cannot identify the violating subset: specifying $W$ requires $\Theta(N)$ bits, while the response transcript reveals only a vanishing amount of information. This recovers the same conclusion as the Fano-based argument.
Interpretation. Under the oracle interaction model, transcript distributions corresponding to different instances are statistically indistinguishable except on exponentially rare events. Consequently, the observable signal distinguishing different instances is extremely weak.
In the near-boundary regime, small perturbations of the parameters may switch the matrix between feasibility and violation while affecting observable quantities only slightly. As a result, transcript distributions induced by feasible and violating instances may remain nearly indistinguishable under restricted interaction, contributing to the resulting information-accessibility barrier.
This suggests that the weakness of the observable signal may not be solely an artifact of the oracle model, but may reflect a property of the instances themselves: the relevant information is globally encoded while only weakly expressed through observable quantities.
Conjecture. There exist distributions over instances of the form $A(u,v)$ under which finding some violating subset $\alpha$ satisfying
@@
can be achieved with non-negligible probability, while reliably recovering a specific violating subset $w^{\star}$ (for example, one satisfying
$\det(A(u,v)_{w^{\star}})\approx-\epsilon$
for a prescribed $\epsilon>0$) remains computationally hard on average.
This gap is not captured by the oracle lower bound established in this work, and suggests a stronger form of information inaccessibility: the solution is globally encoded while the resulting observations remain statistically indistinguishable.
This conjecture suggests that the information-accessibility barrier may extend beyond oracle models, pointing toward a broader class of structured problems in which global encoding and limited observability jointly induce computational difficulty.
@@ 5 Concluding Remarks
A key limitation of the present analysis is the use of a uniform prior and a simplified access model. In realistic settings, such as 3-SAT, clause overlap and adaptive inference introduce complex dependencies that may increase the effective rate of information acquisition. Algorithms such as PPSZ [2] demonstrate that these structural features can be exploited to improve search performance; however, their behavior also suggests that information extraction remains limited and must be accumulated over many steps.
From the information-accessibility perspective, reductions from globally encoded search problems, such as P-matrix violation, to clause-based formulations such as 3-SAT may be viewed as mechanisms that transform weakly accessible global structure into locally exploitable logical structure. The auxiliary variables introduced by such reductions may play a role not only in representation, but also in exposing information in a form accessible to inference procedures such as clause propagation, which can aggregate information across multiple local interactions.
Understanding how structural properties enable higher rates of information extraction, and whether fundamental limits exist within specific interaction models, remains an important direction for future research. More broadly, a central open problem is to characterize the regimes in which information-accessibility barriers arise and to quantify the limits on information flow imposed by problem structure.
A further direction is to study the information-acquisition dynamics of structure-exploiting algorithms such as PPSZ. One may conjecture that such algorithms exhibit distinct phases: an initial phase in which structural features (e.g., critical clauses) enable relatively effective information extraction, followed by a regime in which the remaining uncertainty becomes less accessible, leading to a reduced effective information-acquisition rate. Importantly, this transition need not reflect a loss of structural constraints, such as the existence of critical clauses, but rather a reduction in their accessibility under the algorithmic interaction.
It is natural to ask whether this behavior reflects an intrinsic property of structured search under the given interaction model.